OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

@misty1212 srry for all these questions i had 32 and i needed help w. eight of them and my mom did not understand it so

OpenStudy (misty1212):

HI again!!

OpenStudy (misty1212):

what is hte slope of \(y=-3x+12\) ?

OpenStudy (anonymous):

umm -3

OpenStudy (misty1212):

yes and so will any parallel line have slope \(-3\)

OpenStudy (misty1212):

so either B or D how will you choose?

OpenStudy (anonymous):

how will i

OpenStudy (anonymous):

slope intercept form again

OpenStudy (misty1212):

you can plug in the numbers and see which one works

OpenStudy (misty1212):

or you can write \[y-6=-3(x+1)\]and find \(y\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y-6+6 = -3x -3+6

OpenStudy (anonymous):

sorry did not go step by step kind of rushed through

OpenStudy (misty1212):

ok that works

OpenStudy (misty1212):

you get \[y=-3x+3\] and that should be correct

OpenStudy (anonymous):

ok thank you

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]